OpenStudy (anonymous):

What is a polynomial function in standard form with zeros 1, 2, –3, and –1?

OpenStudy (amoodarya):

y=a(x-1) (x-2) (x-(-3)) (x-(-1))

OpenStudy (anonymous):

ok what do I do after this I'm confused...

OpenStudy (amoodarya):

if standard form is X^n+ax^(n-1)+... so put a=1 then multiply simplify then you have the answer

OpenStudy (amoodarya):

y=(x-1) (x-2) (x-(-3)) (x-(-1)) =(x-1)(x+1)(x-2)(x+3) =(x^2-1)(x^2-x-6)